Electric fields are vector quantities whose magnitudes are measured in units of volts/meter (V/m). Find the resultant electric field when there are two fields, E1 and E2, where E1 is directed vertically upward and has magnitude 98 V/m and E2 is directed 44° to the left of E1 and has magnitude 137 V/m.

magnitude
direction   
Answer #1

here,

let upward direction to be +y direction and left to be in +x direction

E1 = 98 j V/m

the second electric field , E2 = 137 V/m * ( sin(44) i + cos(44) j)

E2 = ( 95.2 i + 98.5 j) V/m

the net electric feild , E = E1 + E2

E = 95.2 i N/C + 196.5 j N/C

the direction of net field , |E| = sqrt(196.5^2 + 95.2^2)

|E| = 218.3 N/C

the direction of field , theta = arctan(95.2/196.5) = 25.8 degree to the left of E1